OpenStudy (anonymous):

Given the parent function of f(x) = x4, what change will occur when the function is changed to -f(2x)? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

graph y = x^4 using desmos

jimthompson5910 (jim_thompson5910):

then add on -x^4 for the second row and finally, add on -(2x)^4 for the third row

jimthompson5910 (jim_thompson5910):

hopefully you see how the graph is changing

OpenStudy (anonymous):

Oh okay thank you so much. I'll let you check my answer later

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Graph opens the opposite way and is narrower?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

What is the end behavior of the function f(x) = -x3 + 2x2 + 4x + 5? Up on the left, up on the right Up on the left, down on the right <----- MY ANSWER Down on the left, up on the right Down on the left, down on the right Is it right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

What are the coordinates of the turning point for the function f(x) = (x + 3)3 + 1? How would I do this one? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the turning point of y = x^3 is (0,0) graph both y = x^3 and y = (x+3)^3 + 1 in desmos and notice how the shifts occur

jimthompson5910 (jim_thompson5910):

hopefully you see the turning point?

OpenStudy (anonymous):

in different rows right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes each row is a separate equation/graph

OpenStudy (anonymous):

it says the turning point for x^3 is (0, 0) then for (x+3)^3 it changes to (-3, 0) but that isn't any of my answer choices.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

now shift things up 1 unit

jimthompson5910 (jim_thompson5910):

you forgot about the +1 at the end

OpenStudy (anonymous):

ohhh so (-3, 1)?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Thank you!! I'll keep doing this but I'll definitely need your help, so it would be great if you just keep an eye out for my questions and help me out if you don't mind :)

OpenStudy (anonymous):

What polynomial identity should be used to prove that 133 = 125 + 8? Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes <----- MY ANSWER @jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 is that right?

jimthompson5910 (jim_thompson5910):

yes because 125 = 5^3 and 8 = 2^3 125 + 8 = 5^3 = 2^3 = sum of cubes

OpenStudy (anonymous):

yep! :D

OpenStudy (anonymous):

Use the graph of the function f(x) = x3 - 3x2 - x + 3 to answer this question: graph of x cubed minus 3 x squared minus x plus 3 What is the average rate of change from x = -1 to x = 2? -3 -1 <----- MY ANSWER 0 3 What about this one?

jimthompson5910 (jim_thompson5910):

it is correct

OpenStudy (anonymous):

Another one i can't seem to figure out it's very confusing @jim_thompson5910

OpenStudy (anonymous):

Select the graph and the description of the end behavior of f(x) = -x3 - 2. graph that decreases from left to right that passes through ordered pairs negative 2, 10; negative 1, 3; 0, 2; 1, 1; and 2, negative 6 This is a cubic function. The end behavior of the graph will decrease to the right and increase to the left. graph that increases from left to right that passes through ordered pairs negative 2, negative 6; negative 1, 1; 0, 2; 1, 3; and 2, 10 This is a cubic function. The end behavior of the graph will increase to the right and decrease to the left. graph that increases from left to right that passes through ordered pairs negative 2, negative 10; negative 1, negative 3; 0, negative 2; 1, negative 1; and 2, 6 This is a cubic function. The end behavior of the graph will increase to the right and decrease to the left. graph that decreases from left to right that passes through ordered pairs negative 2, 6; negative 1, negative 1; 0, negative 2; 1, negative 3; and 2, negative 10 This is a cubic function. The end behavior of the graph will decrease to the right and increase to the left.

jimthompson5910 (jim_thompson5910):

yikes...giant wall of text lol

OpenStudy (anonymous):

LOL sorry I'll try to graph them if u would like?

jimthompson5910 (jim_thompson5910):

it's fine lol, but graphing -x^3 - 2 should help you find the answer

OpenStudy (anonymous):

oh okay let me try that out

jimthompson5910 (jim_thompson5910):

to find ordered pairs on the graph, you can also set up a table

OpenStudy (anonymous):

I got D as the answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how many answer choices are there?

OpenStudy (anonymous):

A through D

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

choice D says graph that decreases from left to right that passes through ordered pairs negative 2, 6; negative 1, negative 1; 0, negative 2; 1, negative 3; and 2, negative 10 This is a cubic function. The end behavior of the graph will decrease to the right and increase to the left. right?

jimthompson5910 (jim_thompson5910):

not sure where it starts

OpenStudy (anonymous):

yeah ur right but I graphed it on desmos and it looked the same as my answer choice @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok let me graph real quick

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

yeah it decreases and it goes through (0,-2)

OpenStudy (anonymous):

alright! another one lol I am extremely sorry

OpenStudy (anonymous):

What are the possible numbers of positive, negative, and complex zeros of f(x) = -3x4 + 5x3 - x2 + 8x + 4? Positive: 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0 Positive: 1; negative: 3 or 1; complex: 2 or 0 Positive: 3 or 1; negative: 1; complex: 2 or 0 Positive: 4 or 2 or 0; negative: 2 or 0; complex: 4 or 2 or 0

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

is it B? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how many sign changes are there in -3x^4+5x^3-x^2+8x+4

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

I count 3 sign changes from -3x^4 to +5x^3 from +5x^3 to -x^2 from -x^2 to +8x

OpenStudy (anonymous):

oh yeah sorry my bad lol ur right

jimthompson5910 (jim_thompson5910):

so there are at most 3 positive real roots

jimthompson5910 (jim_thompson5910):

now we need to find f(-x) f(x) = -3x^4+5x^3-x^2+8x+4 f(-x) = -3(-x)^4+5(-x)^3-(-x)^2+8(-x)+4 f(-x) = -3x^4-5x^3-x^2-8x+4 how many sign changes are there in f(-x) ?

OpenStudy (anonymous):

2? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

just one this time: -8x to +4

jimthompson5910 (jim_thompson5910):

so there is at most 1 negative real root

jimthompson5910 (jim_thompson5910):

now let's say we had 3 positive real roots that leaves 1 root left over, so we'd have 1 negative real root because complex roots always come in pairs

jimthompson5910 (jim_thompson5910):

if we had 2 positive real roots, then we'd have 2 complex roots as well

jimthompson5910 (jim_thompson5910):

one sec, reading something on the descartes rule of signs

OpenStudy (anonymous):

okayy

jimthompson5910 (jim_thompson5910):

ok according to this http://www.purplemath.com/modules/drofsign.htm we subtract by 2

jimthompson5910 (jim_thompson5910):

so we'll have 3 or 1 positive root

OpenStudy (anonymous):

Thanks @jim_thompson5910

OpenStudy (anonymous):

Use the graph of the function f(x) = 3x4 - x3 + 3x2 + x - 3 to answer this question: graph of 3 x to the fourth, minus x cubed, plus 3 x squared, plus x minus 3 What is the average rate of change from x = -1 to x = 0? -6 -3 3 6 I'm stuck on this.. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need to evaluate \[\Large \frac{f(-1)-f(0)}{-1-0}\]

OpenStudy (anonymous):

idk what to do @jim_thompson5910

OpenStudy (anonymous):

Can u tell me what to do? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you know what f(-1) means or represents?

OpenStudy (anonymous):

y?

OpenStudy (anonymous):

is the answer 3? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(-1) is the y value when x = -1

jimthompson5910 (jim_thompson5910):

to find f(-1) you plug x = -1 into the function f(x) and evaluate

jimthompson5910 (jim_thompson5910):

what is f(-1) equal to?

OpenStudy (anonymous):

I got 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(0) = ???

OpenStudy (anonymous):

i thought u said f(-1)??

jimthompson5910 (jim_thompson5910):

yes sorry, I want to know both f(-1) and f(0) f(-1) = 3 is correct now what is f(0) equal to

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

plug x = 0 into the function and evaluate

OpenStudy (anonymous):

i did and i got 0 somehow

jimthompson5910 (jim_thompson5910):

try it again, you should get -3

jimthompson5910 (jim_thompson5910):

-3 isn't the answer btw, it's just the value of f(0)

OpenStudy (anonymous):

oh yeah i got -3 sorry for 3*0^2 I somehow got 3 instead of 0 i wasnt thinking lol

jimthompson5910 (jim_thompson5910):

I see

jimthompson5910 (jim_thompson5910):

f(-1) = 3 f(0) = -3 \[\Large \frac{f(-1)-f(0)}{-1-0} = \frac{3-(-3)}{-1-0} = ???\]

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

-6 is the ultimate answer

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you sooo much!!! how can u make something seem SOOO easy seriously u make me feel dumb all the time LOL @jim_thompson5910